# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

### v2

@@ Abstract
Yuyang Dongthanks: Equal contributionthanks: Current affiliation: SB Intuitions Corp. Email:    yuyang.dong@sbintuitions.co.jp, Nobuhiro Ueda11footnotemark: 1, Krisztián Boros, Daiki Ito, Takuya Sera, Masafumi Oyamada NEC Corporation {dongyuyang,nobuhiro-ueda,krisztian-boros,ito-daiki,takuya-sera,oyamada}@nec.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) lewis2021rag; gao2024ragsurvey; fan2024ragllmsurvey technology enables Large Language Models (LLMs) to provide more accurate responses to user queries by retrieving and leveraging relevant knowledge and documents. These knowledge sources, such as company financial reports, web pages, insurance manuals, and academic papers often contain complex charts, tables, diagrams, and other non-textual elements, collectively referred to as rich documents. Effectively enabling RAG systems to understand and utilize such rich, multimodal content remains a key research challenge.
@@
Therefore, a common challenge is having a VLM process an entire rich document page (text conversion or VQA) at once. One potential solution is to further divide a document page into small regions. Traditional document layout analysis technologies such as DocLayout-YOLO doclayoutyolo can achieve this objective, but they focus on fine-grained analysis, breaking down content into small components such as titles, paragraphs, tables, figures, and captions. This approach could lose important context when processing isolated components and potentially lead to reduced RAG accuracy. In our experiments, conventional layout analysis methods with VLM text conversion and VQA degraded the RAG performance in most cases.
@@ 1.1 Contributions
To address these challenges, we propose SCAN, a novel approach that performs VLM-friendly semantic document layout analysis with “coarse granularity.” Figure 1 compares the result of conventional layout analysis with that of SCAN. SCAN can semantically divide regions into boxes that cover contiguous components related to the same topic. For example, each of the semantic box [3], [4], and [5] corresponds to independent topics of IT Services, Social Infrastructure, and Others.
To train a powerful SCAN model, we annotated more than 24k document pages with semantic layout labels. SCAN model is fine-tuned from pre-trained object detection models with this training data. We also designed post-processing techniques for RAG applications. Figure 2 gives an overview of applying our SCAN model to both textual and visual RAG pipelines. Concretely, each page of an input document is treated as an image and decomposed into semantic chunks by our SCAN model. For textual RAG, the images of semantic chunks are passed to a VLM that performs text conversion, and then input into existing textual RAG systems. For visual RAG, the resulting image chunks can be the retrieval targets that are directly input to existing visual RAG systems.
We evaluate SCAN’s performance using three datasets featuring both English and Japanese documents. Although SCAN is trained on Japanese data, the experiments show that it can achieve good performance on English benchmarks. Experimental results show that in textual RAG, applying SCAN can improve end-to-end performance by 2.7–9.4 points, while in visual RAG, SCAN can enhance end-to-end performance by 5.6–10.4 points. Moreover, although our SCAN requires multiple inferences on VLMs rather than single page-level processing, the total computational time is reduced because of the reduced number of input tokens per inference.
@@ 2.1 Document Layout Analysis
CNN-based models such as DocLayout‐YOLO (doclayoutyolo) and Transformer-based models such as DiT (li2022dit), LayoutLMv3 (huang2022layoutlmv3), and Beehive (docling) have been proposed for high-performance document layout analysis which are trained with synthetic and human-annotated datasets publaynet; docbank; doclaynet2022. Building on these models, several end-to-end document conversion systems have been developed. Docling (docling), Marker (marker), and MinerU (mineru) provide a comprehensive pipeline for document layout analysis and text conversion. In addition, production systems such as Azure Document Intelligence adi and LlamaParse Premium llamaparse are available.
@@ 2.2 VLMs for Document Conversion
Vision-Language Models (VLMs) have emerged as powerful tools for multimodal understanding tasks. Open models such as the Qwen-VL series Qwen2.5-VL; qwen3-vl or InternVL internvl have also demonstrated impressive capabilities in visual document understanding. Moreover, smaller OCR-specialized VLMs, including GOT got, Nougat nougat, DocVLM docvlm, and olmOCR olmocr, have been developed to efficiently handle document text extraction.
@@ 2.3 Textual and Visual RAG
The rapid progress of LLMs has further strengthened RAG, enabling models to inject external knowledge into responses with higher precision and coverage lewis2021rag; gao2024ragsurvey; fan2024ragllmsurvey. Typical RAG pipelines in previous works first converted document images or PDFs into plain text, and only then performed indexing and retrieval over the extracted passages ohr. Recent results show that using VLMs for document text conversion yields better results than traditional OCR tools ohr; orcbenchv2. On the other hand, with the increasing availability of multimodal embedding and VLMs, there is a growing interest in multimodal RAG systems that can directly index and retrieve images, and can leverage VLMs for answer generation visrag; vdocrag; colpali.
@@ 3 Method
Maps, Traveling information 3,383 16,558
@@ 3.1 Dataset Construction
As we are the first to define semantic document layout analysis in this context, we developed our training and evaluation dataset through a rigorous annotation process. We first collected document pages from the Japanese portion of Common Crawl PDFs ccpdf. To ensure diversity in our training data, we first performed clustering on the CC-PDF-JP corpus using image embeddings with the MiniCPM-Visual-Embedding minicpmv model. We then selected balanced number of sample pages from each cluster and included them in our dataset; the statistics for each cluster are provided in Table 1. We finally labeled 24,577 pages and split them into 22,577 pages for training, 1,000 for validation, and 1,000 for testing. Appendix A provides examples of our dataset.
@@
We collaborated with a specialized data annotation company in Tokyo, engaging six expert Japanese annotators, with a reasonable payment and contract. We first conducted a pilot annotation on small number of samples to establish a detailed guideline regarding box granularity, segmentation criteria, and the handling of ambiguous cases. During the pilot stage, we measured inter-annotator agreement using a matching-based IoU (Intersection over Union) metric (described in Section 3.2) and obtained a score above 0.7. Because the IoU metric is relatively sensitive to box misalignment, we considered the agreement sufficiently high. Thus, we adopted a single-annotator-per-sample protocol for the main annotation phase. We also regularly checked annotation quality and updated the guideline when unclear. Appendix B provides details of our annotation guideline.
@@
OHR-Bench
Academic, Administration, Finance,
Law, Manual, Newspaper, Textbook
En 1,261 8,561 8,498
Text, Table, Formula,
Chart, Reading Order
Allganize
Finance, Information Technology, Manufacturing,
Public Sector, Retail/Distribution
Ja 64 2,167 300 Text, Chart, Table
BizMMRAG
Finance, Government, Medical,
Consulting Sectors, Wikipedia
Ja 42 3,924 146 Text, Chart, Table
@@ 3.2 Object Detection Model Fine-tuning
There are two popular families of object detection architectures: the CNN-based (YOLO series yolo1; yolo11) and the Transformer-based (DETR series detr-eccv-2020; rtdetr). We fine-tuned YOLO11-X (56.9M parameters) [^1] https://docs.ultralytics.com/models/yolo11/ and RT-DETR-X (67.4M parameters) [^2] https://docs.ultralytics.com/models/rtdetr/ supported in the Ultralytics framework (under New AGPL-3.0 License) to develop our SCAN models.
To evaluate the fine-tuned models in terms of bounding box granularity and precision, we developed a matching-based IoU (Intersection over Union) metric. This metric first uses the Hungarian algorithm to perform bipartite matching between predicted and ground-truth bounding boxes. And then, it calculates an average IoU over the matched bounding box pairs and unmatched bounding boxes. For unmatched bounding boxes, we assigned an IoU of 0 to penalize excessive or insufficient predictions.
Table 3 shows the two models’ IoU scores on the validation set of our dataset. Because each predicted bounding box has its confidence score, we varied the confidence threshold to select the optimal predicted bounding box set. We also computed the coverage score of the selected bounding boxes, defined as the ratio of the area covered by the selected boxes to the total area of all ground-truth boxes. With the optimal confidence threshold, the IoU score of the RT-DETR-X fine-tuned models is better than that of the YOLO11-X fine-tuned models. RT-DETR-X (confidence: 0.5) achieved the best IoU score of 59.6. However, its coverage score is the lowest (77.7) among all the RT-DETR-X settings. Thus, we selected the model based on RT-DETR-X with a confidence threshold of 0.4 as our primary SCAN model. Appendix D shows examples of the predicted bounding boxes by our SCAN model.
@@ 4.1 Datasets and Settings
We use three datasets to evaluate the RAG performance: one English dataset, OHR-Bench ohr, and two Japanese datasets, Allganize allganize and our in-house BizMMRAG. Each dataset is used for both textual and visual RAG evaluation, which involves answering questions based on retrieved document content. As far as we know, OHR-Bench is the first benchmark to systematically evaluate the cascading impact of OCR on RAG systems. It enables step-by-step evaluation across OCR, retrieval, generation, and overall. Table 4 gives a detailed summary of these datasets. Section 4.1.1, Section 4.1.2, and Appendix E provide more details on the experimental settings.
@@
SCAN RT-DETR-X (86M) Our Dataset (22k)
@@ 4.1.1 Textual RAG Setting
When evaluating English textual RAG performance, we use OHR-Bench ohr and follow the same evaluation protocol. We use BGE-m3 bgem3 and BM25 as retrieval models, and meta-llama/Llama-3.1-8B-Instruct and Qwen/Qwen2-7B-Instruct as answer models to generate answers according to the retrieved top-2 results. We use three metrics: (a) retrieval, which calculates LCS (Longest Common Subsequence) to measure evidence inclusion in retrieved content; (b) generation, which measures the F1-score of QA when provided with the ground truth page; and (c) overall, which calculates the F1-score of QA for the end-to-end RAG pipeline. The F1-score is calculated using precision and recall of common tokens between the generated result and the ground truth. The final scores will be the average score across the combination of two retrieval models and two answer models.
For the Japanese datasets BizMMRAG and Allganize, we employ intfloat/multilingual-e5-large as a retrieval model. For each query, the top-5 retrieved results are fed to the answer generation model, which is GPT-4o (gpt-4o-2024-08-06). To evaluate the generated answers, we adopt the LLM-as-a-judge framework llmasajudge, using GPT-4o to assign an integer score from 1 to 5 to each generated answer. Answers receiving a score greater than 4 are considered correct (assigned a value of 1), while others are considered incorrect (assigned a value of 0). Final accuracy is computed based on these binary scores.
To demonstrate the effectiveness of our SCAN method, we compare several text conversion methods: (1) using VLMs directly for text conversion without any layout analysis; (2) using fine-grained layout analysis methods, including DiT li2022dit, DocLayout-YOLO doclayoutyolo, and Beehive docling, followed by a VLM for text conversion; and (3) our SCAN method followed by a VLM for text conversion. For the VLMs, we use OCR-specialized models, GOT got, Nougat nougat, and olmOCR olmocr, as well as a general VLM, Qwen/Qwen2.5-VL-72B Qwen2.5-VL. For the layout analysis models, we list their architectures and training data in Table 5. Note that our SCAN model’s training data is much smaller than the other two layout analysis models. For the settings (2) and (3), we apply the post-processing described in Section 3.3.
@@ 4.1.2 Visual RAG Setting
We also apply OHR-Bench, BizMMRAG, and Allganize to evaluate visual RAG performance. We use ColQwen2-v1.0 colpali as an image retrieval model with top-5 retrieval and Qwen/Qwen2.5-VL-7B as an answer model.
@@
DiT-Qwen2.5-VL-72B 75.0 54.6 50.0 59.9 (-8.8) 90.1 67.1 62.2 73.2 (++1.5)
DocLayout-YOLO-Qwen2.5-VL-72B 61.7 25.0 28.6 38.4 (-30.3) 49.3 21.1 23.2 31.2 (-40.5)
Beehive-Qwen2.5-VL-72B 70.0 29.6 21.4 40.3 (-28.4) 61.3 40.8 26.8 43.0 (-28.7)
SCAN-Qwen2.5-VL-72B (ours) 81.7 72.7 73.8 76.1 (++7.4) 85.9 85.5 72.0 81.1 (++9.4)
@@ 4.2 Textual RAG Evaluation Results
OHR-Bench. Table 6 presents the comprehensive evaluation results of our SCAN method applied to various VLMs for text conversion in textual RAG. Among conventional approaches, general VLMs demonstrate superior performance, followed by pipeline-based OCR methods, and the last are OCR-specialized small VLMs.
@@
Our SCAN model can improve the performance of VLM-based text conversions. Even the strong baseline of Qwen2.5-VL-72B model achieves an impressive overall score of 31.1% (the ground truth is 36.1%), applying SCAN to Qwen2.5-VL-72B further improves the performance to 33.8%. The performance gains are larger when applying SCAN to OCR-specialized small VLMs. With GOT, our SCAN’s improvement is 6.2 points, enabling this smaller model to achieve competitive performance comparable to much larger VLMs. This finding has important implications for deployment scenarios with computational constraints, suggesting that our semantic layout analysis approach can help bridge the efficiency-performance gap. The generation results exhibit similar improvement patterns. The results also indicate that SCAN’s enhancements for structured content elements such as reading order (RO) and tables (TAB) become increasingly significant. This suggests that the semantic segmentation approach is particularly valuable for preserving the relationships between elements that have spatial dependencies.
@@
On the other hand, applying SCAN slightly degrades the retrieval performance when using Qwen2.5-VL-72B. This is because retrieval is a relatively simpler task within the RAG pipeline, primarily requiring the correct identification of keywords rather than comprehensive document understanding. In contrast, the subsequent question-answering stage demands precise and complete conversion of document content into text, where SCAN’s semantic layout analysis proves particularly advantageous.
@@
We can also see that fine-grained document analysis methods, DocLayout-YOLO doclayoutyolo and Beehive docling, substantially degrade overall performance when used with Qwen2.5-VL-72B. The degradations are particularly severe for structured content types such as tables and charts. These conventional layout analysis methods typically segment documents into small atomic regions, which also break the structure of documents. In contrast, DiT and our semantic box approach improve the strong baseline of Qwen2.5-VL-72B. Although DiT is generally categorized as a conventional fine-grained document layout analysis method, its outputs include coarser segments compared to DocLayout-YOLO and Beehive, demonstrating that relatively coarser segments are more suitable for VLMs. Our SCAN further optimizes this level of granularity: it preserves the integrity of semantically coherent regions, maintains their holistic structure while still providing the organizational benefits of layout analysis. This preservation of semantic unity enables VLMs to process each region with full contextual awareness, resulting in more accurate text conversion and, ultimately, superior RAG performance.
@@
BizMMRAG and Allganize. Table 7 presents the textual RAG performance for Japanese document datasets. We include DiT, DocLayout-YOLO, and Beehive layout analysis models in our experiments, as the performance of layout analysis is not heavily dependent on language. The results have the same trends as the English OHR-Bench evaluation, demonstrating that our SCAN methodology yields substantial improvements over general VLMs. Specifically, on the BizMMRAG dataset, our SCAN-enhanced approach demonstrates notable 7.4 points improvements compared to the baseline Qwen2.5-VL-72B model: text accuracy decreased by 3.3 points, while chart accuracy increased by 20.4 points and table accuracy increased by 4.7 points. We observe similar trends for the Allganize dataset.
We also observe that the performance improvements are larger for Japanese datasets than for the English OHR-Bench. When using Qwen2.5-VL-72B, the gain was 2.7 points on OHR-Bench, but it increases to 7.4 points on BizMMRAG and 9.4 points on Allganize. In general, multilingual VLMs tend to demonstrate relatively higher performance in major languages such as English. Thus, while they can achieve reasonable performance on English documents even when the input images are complex, their performance is likely to degrade substantially on Japanese documents due to greater layout and linguistic complexity. This indicates that SCAN, which mitigates the input image complexity for VLMs, is especially effective for non-major languages such as Japanese, yielding even greater benefits than in English.

### v3

@@ Abstract
Nobuhiro Ueda11footnotemark: 1, Yuyang Dongthanks: Equal contributionthanks: Current affiliation: SB Intuitions Corp. Email:    yuyang.dong@sbintuitions.co.jp, Krisztián Boros, Daiki Ito, Takuya Sera, Masafumi Oyamada NEC Corporation {nobuhiro-ueda,dongyuyang,krisztian-boros,ito-daiki,takuya-sera,oyamada}@nec.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024) technology enables Large Language Models (LLMs) to provide more accurate responses to user queries by retrieving and leveraging relevant knowledge and documents. These knowledge sources, such as company financial reports, web pages, insurance manuals, and academic papers, often contain complex charts, tables, diagrams, and other non-textual elements, collectively referred to as rich documents. Effectively enabling RAG systems to understand and utilize such rich, multimodal content remains a key research challenge.
@@
Therefore, a common challenge is having a VLM process an entire rich document page (text conversion or VQA) at once. One potential solution is to further divide a document page into small regions. Traditional document layout analysis technologies such as DocLayout-YOLO Zhao et al. (2024b) can achieve this objective, but they focus on fine-grained analysis, breaking down content into small components such as titles, paragraphs, tables, figures, and captions. This approach could lose important context when processing isolated components and potentially lead to reduced RAG accuracy. In our experiments, conventional layout analysis methods with VLM text conversion and VQA degraded RAG performance in most cases.
@@ 1.1 Contributions
To address these challenges, we propose SCAN, a novel approach that performs VLM-friendly semantic document layout analysis with “coarse granularity.” Figure 1 compares the result of conventional layout analysis with that of SCAN. SCAN can semantically divide regions into boxes that cover contiguous components related to the same topic. For example, each of the semantic boxes [3], [4], and [5] corresponds to independent topics of IT Services, Social Infrastructure, and Others.
To train a powerful SCAN model, we annotated more than 24k document pages with semantic layout labels. The model is fine-tuned from pre-trained object detection models using this training data. We also designed post-processing techniques for RAG applications. Figure 2 gives an overview of applying our SCAN model to both textual and visual RAG pipelines. Concretely, each page of an input document is treated as an image and decomposed into semantic chunks by our SCAN model. For textual RAG, the images of semantic chunks are passed to a VLM that performs text conversion, and the resulting texts are then input into existing textual RAG systems. For visual RAG, the resulting image chunks can be the retrieval targets that are directly input into existing visual RAG systems.
We evaluate SCAN’s performance using three datasets featuring both English and Japanese documents. Although SCAN is trained on Japanese data, the experiments show that it can achieve good performance on English benchmarks. Experimental results show that in textual RAG, applying SCAN can improve end-to-end performance by 2.7–9.4 points, while in visual RAG, SCAN can enhance end-to-end performance by 5.6–10.4 points. Moreover, although SCAN requires multiple VLM inferences rather than a single page-level inference, the total computational time is reduced because each inference uses fewer input tokens.
@@ 2.1 Document Layout Analysis
CNN-based models such as DocLayout‐YOLO (Zhao et al., 2024b) and Transformer-based models such as DiT (Li et al., 2022), LayoutLMv3 (Huang et al., 2022), and Beehive (Auer et al., 2024) have been proposed for high-performance document layout analysis. These models are trained with synthetic and human-annotated datasets Zhong et al. (2019); Li et al. (2020); Pfitzmann et al. (2022). Building on these models, several end-to-end document conversion systems have been developed. Docling (Auer et al., 2024), Marker (Datalab, 2024), and MinerU (Wang et al., 2024) provide a comprehensive pipeline for document layout analysis and text conversion. In addition, production systems such as Azure Document Intelligence microsoft (2025) and LlamaParse Premium LlamaParse.AI (2025) are available.
@@ 2.2 VLMs for Document Conversion
Vision-Language Models (VLMs) have emerged as powerful tools for multimodal understanding tasks. Open models such as the Qwen-VL series Bai et al. (2025b, a) and InternVL Chen et al. (2024b) have also demonstrated impressive capabilities in visual document understanding. Moreover, smaller OCR-specialized VLMs, including GOT Wei et al. (2024), Nougat Blecher et al. (2024), DocVLM Nacson et al. (2024), and olmOCR Poznanski et al. (2025), have been developed to efficiently handle document text extraction.
@@ 2.3 Textual and Visual RAG
The rapid progress of LLMs has further strengthened RAG, enabling models to inject external knowledge into responses with higher precision and coverage Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024). Typical RAG pipelines in previous works first converted document images or PDFs into plain text, and only then performed indexing and retrieval over the extracted passages Zhang et al. (2025). Recent results show that using VLMs for document text conversion yields better results than traditional OCR tools Zhang et al. (2025); Fu et al. (2024). On the other hand, with the increasing availability of multimodal embeddings and VLMs, there is a growing interest in multimodal RAG systems that can directly index and retrieve images and leverage VLMs for answer generation Yu et al. (2025); Tanaka et al. (2025); Faysse et al. (2024).
@@ 3 Method
Maps, Travel information 3,383 16,558
@@ 3.1 Dataset Construction
As we are the first to define semantic document layout analysis in this context, we developed our training and evaluation dataset through a rigorous annotation process. We first collected document pages from the Japanese portion of the CCpdf corpus Turski et al. (2023). To ensure diversity in our training data, we first performed agglomerative clustering on a small subset of the corpus using image embeddings with the MiniCPM Visual Embedding Rhapsody Group (2024) model. We then selected a balanced number of sample pages close to the centroid of each cluster and included them in our dataset; the statistics for each cluster are provided in Table 1. We finally labeled 24,577 pages and split them into 22,577 pages for training, 1,000 for validation, and 1,000 for testing. Appendix A provides examples of our dataset.
@@
We collaborated with a specialized data annotation company in Tokyo, engaging six expert Japanese annotators under formal contracts with reasonable payment. We first conducted a pilot annotation on a small number of samples to establish a detailed guideline regarding box granularity, segmentation criteria, and the handling of ambiguous cases. During the pilot stage, we measured inter-annotator agreement using a matching-based IoU (Intersection over Union) metric (described in Section 3.2) and obtained a score above 0.7. Because the IoU metric is relatively sensitive to box misalignment, we considered the agreement sufficiently high. Thus, we adopted a single-annotator-per-sample protocol for the main annotation phase. We also regularly checked annotation quality and updated the guideline. Appendix B provides details of our annotation guideline.
@@
OHR-Bench Academic, Administration, Finance, Law, Manual, Newspaper, Textbook En 1,261 8,561 8,498 Text, Table, Formula, Chart, Reading Order
Allganize Finance, Information Technology, Manufacturing, Public Sector, Retail/Distribution Ja 64 2,167 300 Text, Chart, Table
BizMMRAG Finance, Government, Medical, Consulting Sectors, Wikipedia Ja 42 3,924 146 Text, Chart, Table
@@ 3.2 Object Detection Model Fine-tuning
There are two popular families of object detection architectures: the CNN-based (YOLO series Redmon et al. (2016); Khanam and Hussain (2024)) and the Transformer-based (DETR series Carion et al. (2020); Zhao et al. (2024a)). We fine-tuned YOLO11-X [^1] https://docs.ultralytics.com/models/yolo11/ and RT-DETR-X [^2] https://docs.ultralytics.com/models/rtdetr/ supported by the Ultralytics framework (under the AGPL-3.0 license) to develop our SCAN models.
To evaluate the fine-tuned models in terms of bounding box granularity and precision, we developed a matching-based IoU (Intersection over Union) metric. This metric first uses the Hungarian algorithm to perform bipartite matching between predicted and ground-truth bounding boxes. Then, it calculates an average IoU over the matched bounding box pairs and unmatched bounding boxes. For unmatched bounding boxes, we assigned an IoU of 0 to penalize excessive or insufficient predictions.
Table 3 shows the two models’ IoU scores on the validation set of our dataset. Because each predicted bounding box has its confidence score, we varied the confidence threshold to select the optimal predicted bounding box set. We also computed the coverage score of the selected bounding boxes, defined as the ratio of the area covered by the selected boxes to the total area of all ground-truth boxes. With the optimal confidence threshold, the IoU score of the RT-DETR-X fine-tuned model is better than that of the YOLO11-X fine-tuned model. RT-DETR-X (confidence: 0.5) achieved the best IoU score of 59.6. However, its coverage score is the lowest (77.7) among all the RT-DETR-X settings. Thus, we selected the model based on RT-DETR-X with a confidence threshold of 0.4 as our primary SCAN model. Appendix D shows examples of predicted bounding boxes from our SCAN model.
@@ 4.1 Datasets and Settings
We use three datasets to evaluate the RAG performance: one English dataset, OHR-Bench Zhang et al. (2025), and two Japanese datasets, our in-house BizMMRAG and Allganize Allganize.ai (2024). Each dataset is used for both textual and visual RAG evaluation, which involves answering questions based on retrieved document content. As far as we know, OHR-Bench is the first benchmark to systematically evaluate the cascading impact of OCR on RAG systems. It enables step-by-step evaluation across OCR, retrieval, generation, and overall performance. Table 4 gives a detailed summary of these datasets. Section 4.1.1, Section 4.1.2, and Appendix E provide more details on the experimental settings.
@@
SCAN RT-DETR-X (67M) Our Dataset (22k)
@@ 4.1.1 Textual RAG Setting
When evaluating English textual RAG performance, we use OHR-Bench Zhang et al. (2025) and follow the same evaluation protocol. We use BGE-m3 Chen et al. (2024a) and BM25 as retrieval models, and meta-llama/Llama-3.1-8B-Instruct and Qwen/Qwen2-7B-Instruct as answer models to generate answers according to the retrieved top-2 results. We use three metrics: (a) retrieval, which calculates LCS (Longest Common Subsequence) to measure evidence inclusion in retrieved content; (b) generation, which measures the F1-score of QA when provided with the ground truth page; and (c) overall, which calculates the F1-score of QA for the end-to-end RAG pipeline. The F1-score is calculated using precision and recall of common tokens between the generated result and the ground truth. The final scores are the average across the four combinations of two retrieval models and two answer models.
For the Japanese datasets BizMMRAG and Allganize, we employ intfloat/multilingual-e5-large as a retrieval model. For each query, the top-5 retrieved results are fed to the answer generation model, which is GPT-4o (gpt-4o-2024-08-06). To evaluate the generated answers, we adopt the LLM-as-a-judge framework Gu et al. (2025), using GPT-4o to assign an integer score from 1 to 5 to each generated answer. Answers receiving a score greater than 4 are considered correct (assigned a value of 1), while others are considered incorrect (assigned a value of 0). Final accuracy is computed based on these binary scores.
To demonstrate the effectiveness of our SCAN method, we compare several text conversion methods: (1) using VLMs directly for text conversion without any layout analysis; (2) using fine-grained layout analysis methods, including DiT Li et al. (2022), DocLayout-YOLO Zhao et al. (2024b), and Beehive Auer et al. (2024), followed by a VLM for text conversion; and (3) our SCAN method followed by a VLM for text conversion. For the VLMs, we use an OCR-specialized model, GOT Wei et al. (2024), as well as a general VLM, Qwen2.5-VL Bai et al. (2025b). We also include three other models, Nougat Blecher et al. (2024), olmOCR Poznanski et al. (2025), and InternVL2.5 Chen et al. (2024b) for the setting (1) as baselines. For the layout analysis models, we list their architectures and training data in Table 5. Note that our SCAN model’s training data is much smaller than the other three layout analysis models. For the settings (2) and (3), we apply the post-processing described in Section 3.3.
@@ 4.1.2 Visual RAG Setting
We also apply OHR-Bench, BizMMRAG, and Allganize to evaluate visual RAG performance. We use ColQwen2-v1.0 Faysse et al. (2024) as an image retrieval model with top-5 retrieval and Qwen/Qwen2.5-VL-7B as an answer model.
@@
DiT-Qwen2.5-VL-72B 75.0 54.6 50.0 59.9 (-8.9) 90.1 67.1 62.2 73.2 (++1.5)
DocLayout-YOLO-Qwen2.5-VL-72B 61.7 25.0 28.6 38.4 (-30.4) 49.3 21.1 23.2 31.2 (-40.5)
Beehive-Qwen2.5-VL-72B 70.0 29.6 21.4 40.3 (-28.5) 61.3 40.8 26.8 43.0 (-28.7)
SCAN-Qwen2.5-VL-72B (ours) 81.7 72.7 73.8 76.1 (++7.3) 85.9 85.5 72.0 81.1 (++9.4)
@@ 4.2 Textual RAG Evaluation Results
OHR-Bench. Table 6 presents the comprehensive evaluation results of our SCAN method applied to various VLMs for text conversion in textual RAG. Among conventional approaches, VLMs for OCR demonstrate superior performance, followed by OCR-specialized small VLMs.
@@
Our SCAN model can improve the performance of VLM-based text conversions. Despite the strong baseline performance of Qwen2.5-VL-72B, which achieves an impressive overall score of 31.1% (the ground truth is 36.1%), applying SCAN further improves the performance to 33.8%. The performance gains are larger when applying SCAN to OCR-specialized small VLMs. With GOT, our SCAN’s improvement is 6.2 points, enabling this smaller model to achieve competitive performance comparable to much larger VLMs. This finding has important implications for deployment scenarios with computational constraints, suggesting that our semantic layout analysis approach can help bridge the efficiency-performance gap. The generation results exhibit similar improvement patterns. The results also indicate that SCAN’s enhancements for structured content elements such as reading order (RO) and tables (TAB) become increasingly significant. This suggests that the semantic segmentation approach is particularly valuable for preserving the relationships between elements that have spatial dependencies.
@@
On the other hand, applying SCAN slightly degrades the retrieval performance when using Qwen2.5-VL-72B. This is because retrieval is a relatively simple task within the RAG pipeline, primarily requiring the correct identification of keywords rather than comprehensive document understanding. In contrast, the subsequent question-answering stage demands precise and complete conversion of document content into text, where SCAN’s semantic layout analysis proves particularly advantageous.
@@
We can also see that fine-grained document analysis methods, DocLayout-YOLO Zhao et al. (2024b) and Beehive Auer et al. (2024), substantially degrade overall performance when used with Qwen2.5-VL-72B. The degradations are particularly severe for structured content types such as tables and charts. These conventional layout analysis methods typically segment documents into small atomic regions, which also break the structure of documents. In contrast, DiT and our semantic box approach improve the strong baseline of Qwen2.5-VL-72B. Although DiT is generally categorized as a conventional fine-grained document layout analysis method, its outputs include coarser segments compared to DocLayout-YOLO and Beehive, demonstrating that relatively coarser segments are more suitable for VLMs. Our SCAN further optimizes this level of granularity: it preserves the integrity of semantically coherent regions, maintains their holistic structure while still providing the organizational benefits of layout analysis. This preservation of semantic unity enables VLMs to process each region with full contextual awareness, resulting in more accurate text conversion and, ultimately, superior RAG performance.
@@
BizMMRAG and Allganize. Table 7 presents the textual RAG performance for Japanese document datasets. We include DiT, DocLayout-YOLO, and Beehive layout analysis models in our experiments, as the performance of layout analysis is not heavily dependent on language. The results have the same trends as the English OHR-Bench evaluation, demonstrating that our SCAN methodology yields substantial improvements over a capable VLM. Specifically, on the BizMMRAG dataset, our SCAN-enhanced approach demonstrates a notable 7.3-point improvement compared to the baseline Qwen2.5-VL-72B model: text accuracy decreased by 3.3 points, while chart accuracy increased by 20.4 points and table accuracy increased by 4.7 points. We observe similar trends for the Allganize dataset.
We also observe that the performance improvements are larger for Japanese datasets than for the English OHR-Bench. When using Qwen2.5-VL-72B, the gain was 2.7 points on OHR-Bench, but it increases to 7.3 points on BizMMRAG and 9.4 points on Allganize. In general, multilingual VLMs tend to demonstrate relatively higher performance in major languages such as English. Thus, while they can achieve reasonable performance on English documents even when the input images are complex, their performance is likely to degrade substantially on Japanese documents due to greater layout and linguistic complexity. This indicates that SCAN, which mitigates the input image complexity for VLMs, is especially effective for non-major languages such as Japanese, yielding even greater benefits than in English.